At 22Degree C an excess amount of a generic metal

Please show all work!

Ch 13B, 7
Answer #1

M(OH)2 ---> M+2 + 2OH-

From the stoichiometry of the above eqaution we find that for every mol of M, 2 mols of OH is consumed. Therefore, to find the Ksp,

Ksp=[M2+][OH-]. Instead of using the [M2+], we can substitute in [OH-/2],

First pH = 10.16

So pOH =14-10.16=3.84


pOH=3.84

pOH = 3.84
[OH-] = 10^-3.84
[OH-] = 1.45*10^-4M


[M+2] = 1.45*10^-4M/2=7.25*10^-5M


ksp= [M+2]*[OH-]^2
= [7.25*10^-5]*[ 1.45*10^-4]^2

=1.52*10^-12